# Model-Based Diffusion Sampling for Predictive Control in Offline Decision Making

### v1

@@ 3.2 Alternating Diffusion Sampling
To generate trajectories, we employ an alternating diffusion sampling scheme (Algorithm 1) that decomposes denoising into two complementary updates: one enforcing dynamics feasibility and the other promoting task alignment. Starting from Gaussian noise over states and actions, each reverse step first applies the dynamics model to refine states conditioned on the current actions, projecting them toward the manifold of feasible transitions, followed by the planner, which jointly denoises states and actions to restore task structure and dataset consistency.
Classifier-free guidance is applied in both steps, blending conditional and unconditional predictions to control the strength of task conditioning. During the reverse process, the planner pushes samples toward high-reward, task-aligned regions, while the dynamics model counteracts drift and enforces feasibility. The alternating composition thus functions like a dialogue: the planner expands trajectories toward task objectives, and the dynamics model regularizes them to respect system transitions.
1:Planner model $\epsilon^{\mathrm{pl}}_{\theta}(\tau_{x}^{k},\tau_{u}^{k}\mid k,x_{0},y)$; Dynamics model $\epsilon^{\mathrm{dyn}}_{\vartheta}(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},y)$; guidance scale $\omega$; condition $y$; initial state $x_{0}$, temperature $\alpha$
2:Initialize actions $\tau_{u}^{k}\sim\mathcal{N}(0,\alpha I)$, states $\tau_{x}^{k}\sim\mathcal{N}(0,\alpha I)$
3:for $k=K,\dots,1$ do
[…]
5:  $\hat{\epsilon}_{x}\leftarrow\;\omega\epsilon_{\vartheta}^{\mathrm{dyn}}(\tau^{k}_{x}|\tau^{k}_{u},k,x_{0},y)+(1-\omega)\epsilon_{\vartheta}^{\mathrm{dyn}}(\tau^{k}_{x}|\tau^{k}_{u},k,x_{0},\emptyset)$
[…]
8:  $\hat{\epsilon}_{\tau}\leftarrow\;\omega\epsilon_{\theta}^{\mathrm{pl}}(\tilde{\tau}_{x},\tau_{u}^{k}\mid k,x_{0},y)+(1-\omega)\epsilon_{\theta}^{\mathrm{pl}}(\tilde{\tau}_{x},\tau_{u}^{k}\mid k,x_{0},\emptyset)$
9:  $(\mu^{k-1}_{\tau},\Sigma^{k-1}_{\tau})\leftarrow\mathrm{Denoise}(\tilde{\tau}_{x},\tau_{u}^{k},\hat{\epsilon}_{\tau})$
10:  $\tau_{x}^{k-1},\tau_{u}^{k-1}\sim\mathcal{N}(\mu^{k-1}_{\tau},\alpha\Sigma^{k-1}_{\tau})$
11:end for
12:return $\tau=(\tau_{x}^{0},\tau_{u}^{0})$
@@ 3.3 Constrained Control with MPDiffuser
1:initial state $x_{0}$ , num. candidates NN, remaining budget $B_{\mathrm{rem}}$, return scale $\lambda_{R}$, cost scale $\lambda_{C}$, reward model $\hat{r}$, cost model $\hat{c}$, discount factor $\gamma$
@@
3:for $i=1$ to NN do
4:  $\hat{J}_{i}\leftarrow\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(x_{t}^{(i)},u_{t}^{(i)})$,  $\hat{C}_{i}\leftarrow\sum_{t=0}^{H-1}\gamma^{t}\hat{c}(x_{t}^{(i)},u_{t}^{(i)})$
@@
8:  return $\tau^{\arg\max_{i\in\mathcal{F}}\hat{J}_{i}}$ $\triangleright$ pick highest reward among feasible
@@
10:  return $\tau^{\arg\min_{i}\hat{C}_{i}}$ $\triangleright$ fallback: pick least-cost if none feasible
@@ 3.4 Rationale behind the algorithm development
Directly characterizing $q^{\ast}$ is not possible, as it is an abstract construction combining $p_{\mathrm{pl}}$ and $p_{\mathrm{dyn}}$, and we do not have samples from it to fit a diffusion model. Nevertheless, sampling from $q^{\ast}$ can in principle be achieved via its score function $s_{q^{\ast}}$, which determines the probability–flow dynamics. The exact score is intractable, but by analogy with classifier guidance and related methods we approximate it as a sum of individual scores:
@@
where $s_{q^{\ast}}$, $s_{p^{\mathrm{pl}}}$, $s_{p^{\mathrm{dyn}}}$ denotes the score function of the corresponding probabilities. A natural approach is to update trajectories jointly with this combined score, but in practice such updates can be unstable because planner and dynamics gradients often differ in scale or curvature (cf. Appendix I). Our algorithm instead alternates between planner and dynamics updates. This design is motivated by operator-splitting methods (hairer2006geometric; trotter1959product), which approximate the flow of a combined system by alternating short steps under each component. Although both models share the same architecture and training data, the dynamics model focuses exclusively on state prediction, while the planner models both states and actions. This specialization allows the dynamics model to capture transition structure more accurately, yielding a stronger state-consistency signal. Alternating the two thus combines the planner’s task alignment with the dynamics model’s precision, guiding sampling effectively toward $q^{\ast}$.
@@ 4 Experiments
We evaluate our method across diverse settings to demonstrate its effectiveness, versatility, and practicality. Our experimental evaluation includes: (1) offline decision making on D4RL benchmark tasks, including adaptation to novel dynamics, assesment of feasibility of the generated sequences and leveraging random data for dynamics learning (Sec.4.1); (2) constrained offline decision making on safety-critical DSRL benchmarks with cost constraints, and a study on classical Pendulum environment highlighting significance of dynamic feasibility for trajectory ranking (Sec. 4.2); (3) a preliminary study extending our framework to handle visual inputs (Sec. 4.3); (4) real-world deployment on a Unitree Go2 quadruped robot to demonstrate the practicality of MPDiffuser (Sec. 4.4); (4) a linear control system with stochastic expert data to validate the approach on a well-understood theoretical setup (App. E); (5) additional ablations, including: the mode of initial-state conditioning (App. D), evaluating robustness to modeling errors in our dynamics model (App. F), using two models versus a single one with more diffusion steps (App. H), alternating versus combined score updates (App. I), and the role of conditioning in the dynamics model (App. J), the impact of causal architectures (App. K), sensitivity to guidance scale (App. M) and number of samples for ranking, performance degradation due to distribution shift between components (App. L).Code and experiments are available at: https://anonymous.4open.science/status/MPD-Submission-126B
@@ 4.1 Offline Decision Making
Results on Standardized Benchmarks (D4RL): We evaluate MPDiffuser, in two configurations: (i) MPDiffuser using a single trajectory sample, and (ii) MPDiffuser +Rank using multiple samples (64), where the ranker selects highest return trajectory among sampled candidates using a learned reward model. Experiments are conducted on the D4RL benchmark (d4rl). We compare against standard baselines such as Behavior Cloning (BC) and Decision Transformer (DT) (chen2021decision), a model-based offline RL algorithm (COMBO) (yu2021combo), as well as recent diffusion-based methods including IDQL (hansen2023idql), Diffusion MPC (D-MPC) (zhou2025diffusion), Decision Diffuser (ajay2023conditional) and Diffuser (janner2022planning). To isolate the effect of the alternating sampling scheme, we include results using the planner alone.
@@
Dataset Environment BC DT COMBO IDQL Diffuser
Decision
Diffuser
D-MPC Planner MPDiffuser MPDiffuser+Rank
Med-Exp
Hopper $52.5$ $107.6$ 111.1 105.3 $107.2$ $\mathbf{111.8}$ $109.5$ $09.9\pm 1.1 11$$10.4\pm 0.0$
@@
Medium
Hopper $52.9$ $67.6$ 97.2 63.1 $58.5$ $79.3$ $61.2$ $97.6$ $97.9\pm 0.3$ $\mathbf{98.4}\pm 0.4$
@@
Med-Replay
Hopper $18.1$ $82.7$ 89.5 82.4 $96.8$ $\mathbf{100.0}$ $92.5$ $92.1$ $98.2\pm 0.3$ $98.3\pm 0.7$
@@
Table 1: Performance on D4RL benchmark tasks. We report normalized average scores with corresponding standard deviations under the standard D4RL evaluation protocol (d4rl). Results are averaged over 5 independent runs, each evaluated on 50 rollouts. Overall, MPDiffuser outperforms prior baselines, while MPDiffuser +Rank provides further improvements by selecting higher-quality trajectories.
@@
Num. random traj. 0 2000 4000 6000 8000 10000
MPDiffuser 0.75 0.81 0.79 0.78 0.82 0.86
D-MPC 0.60 0.68 0.56 0.53 0.55 0.60
Table 2: MPDiffuser can harness suboptimal data. Success rate on FetchPickAndPlace when training the dynamics model with additional random trajectories, while keeping the planner fixed to 1,000 expert demonstrations. MPDiffuser’s performance improves with more random data, showing that inexpensive, trajectories can enhance learning, whereas D-MPC shows no consistent improvement.
Figure 4: Dynamics consistency of sampled trajectories. Mean state error over the prediction horizon for: block position in world coordinates, block position relative to the end-effector, and all state dimensions combined. MPDiffuser achieves lower state prediction error compared to Decision Diffuser and the planner-only baseline, indicating improved consistency with system dynamics.
Leveraging Random Data for Dynamics Learning: We consider FetchPickAndPlace environment, where a robotic arm must bring a block to a target location. Both models are conditioned on the goal, and ranker picks trajectories by minimum block–goal distance, with success defined as bringing block close enough to the goal position. The planner is trained on $1000$ expert demonstrations, while the dynamics model additionally uses trajectories obtained by applying random actions.
Table 2 reports success rates. Adding random trajectories for dynamics training improves performance, even though the planner relies solely on expert data. This demonstrates a key benefit of our compositional framework: while training planners require high-quality data, the dynamics model can effectively exploit inexpensive random data to enhance feasibility and performance. Additionally, we evaluate D-MPC under the same setting. For a fair comparison, we train its planner and dynamics models following the original formulation (zhou2025diffusion), using architectures comparable to ours. We find that adding more data does not improve D-MPC’s performance, as its dynamics model is applied only after planning for candidate filtering and does not influence action proposals. Consequently, a better dynamics model merely refines selection rather than generation. In contrast, MPDiffuser incorporates the dynamics model directly into the sampling process, allowing its improvements to immediately enhance the quality of generated trajectories.
Assessing Dynamics Consistency of Sampled Trajectories: We evaluate the dynamics consistency of trajectories generated by different diffusion models on the FetchPickAndPlace environment. Each model is trained on 1,000 expert demonstrations. For evaluation, we sample 250 random initial states, generate trajectories using each model, and compare the simulated rollouts (obtained by executing the generated actions) with the corresponding diffused state trajectories. The mean errors are reported in Fig. 4. MPDiffuser demonstrates stronger dynamics consistency than both Decision Diffuser and the planner-only baseline, while D-MPC achieves a comparable state error. However, despite the similar state deviation, MPDiffuser achieves a notably higher success rate (75%) than D-MPC (60%) as noted in Table 2, highlighting that our alternating planner–dynamics sampling more effectively balances task fidelity with dynamic feasibility.
Adapting to Novel Dynamics: To assess our method’s adaptability to changing system dynamics, we follow the experimental protocol from zhou2025diffusion. Accordingly, we train models on the D4RL walker2d-medium dataset and simulate a hardware defect by limiting the torque of one ankle joint to the range [0.5,0.5][-0.5,0.5]. Table 6 summarizes the results. Originally, all methods achieve similar performance. When deployed directly under the defect, however, both Diffuser and D-MPC suffer substantial performance drops, while MPDiffuser maintains significantly higher returns.
@@
Figure 5: Walker2D visualization (highlights defective joint)
Diffuser D-MPC Planner MPD MPD+Rank
Original 79.6 76.2 75.9 77.6 77.6
Pre-FT w/ defect 25.9 22.7 58.6 58.6 51.0
Post-FT w/ defect 6.8 30.7 56.0 66.4 63.4
Figure 6: MPDiffuser can adapt to novel dynamics. Performance before and after fine-tuning (FT) under defect. Diffuser and D-MPC results from zhou2025diffusion. Only the dynamics model is tuned for MPDiffuser and the ranker for MPDiffuser+Rank.

### v2

@@ 3.2 Alternating Diffusion Sampling
To generate trajectories, we employ an alternating diffusion sampling scheme (Alg. 1) that decomposes denoising into two complementary updates: one enforcing dynamics feasibility and the other promoting task alignment. Starting from Gaussian noise, each reverse step first applies the dynamics model to refine states conditioned on the current actions, projecting them toward the manifold of feasible transitions, followed by the planner, which jointly denoises states and actions to restore task structure and dataset consistency.
Classifier-free guidance is applied in both steps, to encourage task conditioning. During the reverse process, the planner pushes samples toward task-aligned regions, while the dynamics model counteracts drift and enforces feasibility. The alternating composition thus functions like a dialogue: the planner expands trajectories toward task objectives, and the dynamics model regularizes them to system transitions.
@@ 3.3 Constrained Control with MPDiffuser
0: Planner $\epsilon^{\mathrm{pl}}_{\theta}(\tau_{x}^{k},\tau_{u}^{k}\mid k,x_{0},y)$; Dynamics $\epsilon^{\mathrm{dyn}}_{\vartheta}(\tau_{x}^{k}\mid\tau_{u}^{k},k,x_{0},y)$; guidance scale $\omega$; condition $y$; initial state $x_{0}$, temperature $\alpha$
1: Initialize trajectory $\tau_{x}^{k},\tau_{u}^{k}\sim\mathcal{N}(0,\alpha I)$
2:for $k=K,\dots,1$ do
2:   $\triangleright$ Dynamics step: update states only
3:  $\hat{\epsilon}_{x}\leftarrow\mathrm{CFG}_{\omega}\!\left(\epsilon_{\vartheta}^{\mathrm{dyn}}(\tau^{k}_{x}|\tau^{k}_{u},k,x_{0},y)\right)$ (eq. (1))
[…]
4:   $\triangleright$ Planner step: update states & actions jointly
5:  $\hat{\epsilon}_{\tau}\leftarrow\mathrm{CFG}_{\omega}\!\left(\epsilon_{\theta}^{\mathrm{pl}}(\tilde{\tau}_{x},\tau_{u}^{k}\mid k,x_{0},y)\right)$ (eq. (1))
6:  $(\mu^{k-1}_{\tau},\Sigma^{k-1}_{\tau})\leftarrow\mathrm{Denoise}(\tilde{\tau}_{x},\tau_{u}^{k},\hat{\epsilon}_{\tau})$
7:  $\tau_{x}^{k-1},\tau_{u}^{k-1}\sim\mathcal{N}(\mu^{k-1}_{\tau},\alpha\Sigma^{k-1}_{\tau})$
8:end for
9:return $\tau=(\tau_{x}^{0},\tau_{u}^{0})$
@@
0: initial state $x_{0}$ , num. candidates NN, condition $y$, remaining budget $b_{\mathrm{rem}}$, reward model $\hat{r}$, cost model $\hat{c}$, discount factor $\gamma$
@@
2:for $i=1$ to NN do
3:  $\hat{J}_{i}\leftarrow\sum_{t=0}^{H-1}\gamma^{t}\hat{r}(x_{t}^{(i)},u_{t}^{(i)})$
4:  $\hat{C}_{i}\leftarrow\sum_{t=0}^{H-1}\gamma^{t}\hat{c}(x_{t}^{(i)},u_{t}^{(i)})$
@@
8:  return highest return feasible trajectory
@@
10:  return minimum cost trajectory
@@ 3.4 Rationale behind the algorithm development
Directly characterizing $q^{\ast}$ is not possible, as it is an abstract construction combining $p_{\mathrm{pl}}$ and $p_{\mathrm{dyn}}$, and we do not have samples from it to fit a diffusion model. Nevertheless, sampling from $q^{\ast}$ can in principle be achieved via its score function $s_{q^{\ast}}$, which determines the probability–flow dynamics. The exact score is intractable, but by analogy with classifier guidance we approximate it as sum of scores:
@@
where $s_{q^{\ast}}$, $s_{p^{\mathrm{pl}}}$, $s_{p^{\mathrm{dyn}}}$ denotes the score function of the corresponding distributions. A natural approach is to sample using this combined score, but in practice such updates can be unstable because planner and dynamics gradients often differ in scale or curvature (cf. Appendix I). Our algorithm instead alternates between planner and dynamics updates. This design is motivated by operator-splitting (Hairer et al., 2006; Trotter, 1959), which approximate the flow of a combined system by alternating short steps under each component. Although both models share the same architecture and training data, the dynamics model focuses exclusively on state prediction, while the planner models both states and actions. This allows the dynamics model to capture transition structure more accurately, yielding a stronger state-consistency signal. Thus, alternating sampling combines the planner’s task alignment with the dynamics model’s precision, guiding sampling effectively toward $q^{\ast}$.
@@ 4 Experiments
We evaluate our method across diverse settings to demonstrate its effectiveness, versatility, and practicality. Our experimental evaluation includes: (1) offline decision making on D4RL benchmark tasks, including adaptation to novel dynamics, assesment of feasibility of the generated sequences and leveraging random data for dynamics learning (Sec.4.1); (2) constrained offline decision making on safety-critical DSRL benchmarks with cost constraints, and a study on Pendulum environment highlighting importance of dynamic feasibility for ranking (Sec. 4.2); (3) a preliminary study extending our framework to handle visual inputs (Sec. 4.3); (4) real-world deployment on a Unitree Go2 quadruped robot to demonstrate the practicality of MPDiffuser (Sec. 4.4).
Additional studies are deferred to the appendix, including: (i) a linear control system for validation in a well-understood theoretical setting (App. E); (ii) ablations on initial-state conditioning (App. D), robustness to dynamics modeling errors (App. F), planner–dynamics alternation versus additional diffusion steps (App. H), alternating versus combined score updates (App. I), conditioning in the dynamics model (App. J), the impact of causal architectures (App. K), sensitivity to guidance scale and number of samples (App. M), and performance degradation under distribution shift between components (App. L). Code to reproduce our results is available at https://anonymous.4open.science/r/MPD-Submission-126B.
@@ 4.1 Offline Decision Making
Results on Standardized Benchmarks (D4RL): We evaluate MPDiffuser , in two configurations: (i) MPDiffuser using a single trajectory sample, and (ii) MPDiffuser +Rank using multiple samples (64), where the ranker selects highest return trajectory among sampled candidates using a learned reward model. Experiments are conducted on the D4RL benchmark (Fu et al., 2021). We compare against standard baselines such as Behavior Cloning (BC) and Decision Transformer (DT) (Chen et al., 2021), a model-based offline RL algorithm (COMBO) (Yu et al., 2021b), as well as recent diffusion-based methods including IDQL (Hansen-Estruch et al., 2023), Diffusion MPC (D-MPC) (Zhou et al., 2025), Decision Diffuser (Ajay et al., 2023) and Diffuser (Janner et al., 2022). To isolate the effect of the alternating sampling scheme, we include results using the planner alone.
@@
Dataset Environment BC DT COMBO IDQL Diffuser Decision Diffuser D-MPC Planner MPDiffuser MPDiffuser+Rank
Med-Exp Hopper $52.5$ $107.6$ 111.1 105.3 $107.2$ $\mathbf{111.8}$ $109.5$ $09.9\pm 1.1 11$$10.4\pm 0.0$
@@
Medium Hopper $52.9$ $67.6$ 97.2 63.1 $58.5$ $79.3$ $61.2$ $97.6$ $97.9\pm 0.3$ $\mathbf{98.4}\pm 0.4$
@@
Med-Replay Hopper $18.1$ $82.7$ 89.5 82.4 $96.8$ $\mathbf{100.0}$ $92.5$ $92.1$ $98.2\pm 0.3$ $98.3\pm 0.7$
@@
Table 1: Performance on D4RL benchmark tasks. We report normalized average scores with corresponding standard deviations under the standard D4RL evaluation protocol (Fu et al., 2021). Results are averaged over 5 independent runs, each evaluated on 50 rollouts. MPDiffuser outperforms prior baselines, while MPDiffuser+Rank provides further improvements by selecting higher-quality trajectories.
Leveraging Random Data for Dynamics Learning: We consider FetchPickAndPlace task, where a robot arm must bring a block to a target location, with success defined as bringing block close enough to the goal position. Models are conditioned on the goal, and ranker picks trajectories by minimum block–goal distance. The planner is trained on $1000$ expert demonstrations, while the dynamics model uses additional trajectories obtained by applying random actions.
Table 2 reports success rates. Adding random trajectories for dynamics training improves performance, even though the planner relies solely on expert data. This demonstrates a key benefit of our compositional framework: while training planners require high-quality data, the dynamics model can effectively exploit inexpensive random data to enhance feasibility and performance. Additionally, we evaluate D-MPC under the same setting. For a fair comparison, we train its planner and dynamics models following the original formulation (Zhou et al., 2025), using architectures comparable to ours. We find that adding more data does not improve D-MPC’s performance, as its dynamics model is applied only after planning for candidate filtering and does not influence action proposals. Consequently, a better dynamics model merely refines selection rather than generation. In contrast, MPDiffuser incorporates the dynamics model directly into the sampling process, allowing its improvements to immediately enhance the quality of generated trajectories.
Assessing Dynamics Consistency of Sampled Trajectories: We evaluate the dynamics consistency of trajectories generated by different diffusion models on the FetchPickAndPlace environment. Each model is trained on $1,000$ expert demonstrations. For evaluation, we sample $250$ random initial states, generate trajectories using each model, and compare the simulated rollouts (obtained by executing the sampled actions) with the diffused state trajectories. The mean errors are reported in Fig. 4. MPDiffuser demonstrates stronger dynamics consistency than both Decision Diffuser and the planner-only baseline, while D-MPC achieves a comparable performance. However, despite the similar state deviation, MPDiffuser achieves a notably higher success rate (75%) than D-MPC (60%) as noted in Table 2, highlighting that our alternating sampling scheme balances task fidelity with dynamic feasibility more effectively. This difference arises because MPDiffuser integrates the dynamics model directly into the sampling process, actively correcting trajectories at each diffusion step, whereas D-MPC applies dynamics only as a post-hoc filtering.
Adapting to Novel Dynamics: To assess our method’s adaptability to changing system dynamics, we follow the experimental protocol from Zhou et al. (2025). Accordingly, we train models on the D4RL walker2d-medium dataset and simulate a hardware defect by limiting the torque of one ankle joint to the range [0.5,0.5][-0.5,0.5]. Table 3 summarizes the results, for Diffuser and D-MPC results are obtained from Zhou et al. (2025). Originally, all methods achieve similar returns; yet, when deployed under the defect, both Diffuser and D-MPC suffer substantial performance drops, while MPDiffuser maintains significantly higher returns.
@@
Num. rand. traj. MPD D-MPC
0 0.75 0.60
2000 0.81 0.68
4000 0.79 0.56
6000 0.78 0.53
8000 0.82 0.55
10000 0.86 0.60
Table 2: MPDiffuser can harness suboptimal data. Success rate versus number of random trajectories for dynamics training.
Figure 4: Dynamics consistency of sampled trajectories. Mean state error over the prediction horizon for block position and all state dimensions.
Figure 5: Walker2D illustration, higlights defective joint
Original Pre-FT Post-FT
Diffuser 79.6 25.9 6.8
D-MPC 76.2 22.7 30.7
Planner 75.9 58.6 56.0
MPD 77.6 58.6 66.4
MPD+Rank 77.6 51.0 63.4
Table 3: MPDiffuser can adapt to novel dynamics. Performance before and after fine-tuning (FT) under defect.
BC-All COptiDICE BC-Safe CPQ BCQ-Lag CDT MPDiffuser
Return Cost Return Cost Return Cost Return Cost Return Cost Return Cost Return Cost
HopperVelocity 0.65 6.39 0.13 1.51 0.36 0.67 0.14 2.11 0.78 5.02 0.63 0.61 0.81 0.37
Walker2dVelocity 0.79 3.88 0.12 0.74 0.79 0.04 0.04 0.21 0.79 0.17 0.78 0.06 0.80 0.27
HalfCheetahVelocity 0.97 13.1 0.65 0.0 0.88 0.54 0.29 0.74 1.05 18.21 1.0 0.01 0.98 0.77
PointGoal1 0.65 0.95 0.49 1.66 0.43 0.54 0.57 0.35 0.71 0.98 0.69 1.12 0.74 0.88
PointCircle1 0.79 3.98 0.86 5.51 0.41 0.16 0.43 0.75 0.54 2.38 0.59 0.69 0.58 0.94
CarGoal1 0.39 0.33 0.35 0.54 0.24 0.28 0.79 1.42 0.47 0.78 0.66 1.21 0.63 0.92
CarCircle1 0.72 4.39 0.70 5.72 0.37 1.38 0.02 2.29 0.73 5.25 0.60 1.73 0.50 0.85
Table 4: Performance on DSRL benchmark tasks. Normalized returns and costs on the DSRL benchmark, with gray entries indicating unsafe behavior. MPDiffuser achieves competitive returns while maintaining safety, demonstrating effective safety-performance balance.